# Combine and conquer: model averaging for out-of-distribution forecasting

### v2

@@ 3.4 Model averaging: specification and estimation
Similar to the MLP models implemented for the training on the data sets (see Section 3.3), we implemented the MLP for the model averaging in PyTorch Paszke et al. (2019) too. This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in e.g. SciKit-learn. As explanatory features, we included three distance-based variables: the raw trip distance, its logarithmic transformation, and its squared term. While in theory the linear distance alone should suffice, we found that adding the log- and squared transformations provided a small but consistent improvement in model performance. To tune the hyperparameters, we performed a grid search using a simple 75%–25% train–test split. Ideally, a dedicated validation set would have been used for this purpose—separate from the data employed in training the individual models, see Hastie (2009), who argues that not doing so might result in the meta-model assigning unfairly high weights to models with higher complexity. However, we deliberately refrained from further partitioning the data to avoid diminishing the already limited estimation sample and to ensure that the test set remained completely untouched for the final performance evaluation. The following parameter spaces are used: batch size (64, 128, 256), number of hidden layers and nodes (two or three hidden layers with each 10, 20, 40, or 60 nodes), learning rate (0.001, 0.005, 0.01) and L2 regularisation (0, 0.0001, 0.001). For the DECISIONS data set, the best-performing model has two hidden layers with 20 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.001. For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
@@
We repeated the training 100 times using the optimal configuration. As we aim for high prediction performance, we retained only the top 20 per cent of models [^2] This selection was based on the log-likelihood. For this, we used the full data, i.e. training and validation data, but excluded the test data., discarding those that became trapped in poor local maxima. This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
Figure 2 shows the structure for our model averaging approach. As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model. The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model-averaging MLP.
@@ 3.5.1 Performance of individual sub-models
Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 3 apply in each of the 6 distance segments used in estimation (i.e. segments 3-8). Across the distance segments, there are some where XGB outperforms MLP in the DECISIONS data, or MLP outperforms XGB in LPMC (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases. These findings already show that different models perform differently well across trip distances. Interestingly, while the LPMC results show that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
@@
We next turn to the test results, shown in the right two graphs in Figure 3. We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data. For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for all models, with an average gain in per observation LL of $0.096$ for NL, $0.095$ for DFT, $0.062$ for MLP, and $0.054$ for XGB. The same happens for LPMC models, the same is observed, with improvements by $0.12$ for NL and DFT, $0.11$ for MLP and $0.094$ for XGB. Overall, these results do not suggest any overfitting, at least within the distribution of distances covered by the estimation data.
The findings for segments 2 and 9 on the test data largely confirm those we saw in the two graphs on the left in Figure 3, except for MLP and XGB performing more poorly in the test data for segment 9. We finally turn to segments 1 and 10. From a behavioural perspective, we see that for the DECISIONS data, mode choice on very short tips is much easier to predict. For short trips in the LPMC dataset, there is a clear advantage for the behavioural models. For the longest trips (segment 10), the trend is less clear. In the DECISIONS data, MLP performs comparatively well. In the LPMC data, both data-driven models do comparatively well, while MNL suffers a big drop.
@@ 3.5.2 Model averaging results
For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e. distances on which the individual sub-models were estimated. In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP for DECISIONS, while it also increases for LPMC. These findings are in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
Once we go outside the $D_{train,sub}$ interval, the pattern changes. In both datasets, we observe that for trips with distances below $d_{a,sub}$ (i.e. segments 1&2), the weight for behavioural models increases. In the DECISIONS data, only a small share goes to NL, with quite equal weight for the other three models. On the other hand, in LPMC, MNL essentially obtains all the weight. For longer trips, i.e. above $d_{b,sub}$, NL gets the vast majority of the weight in the DECISIONS data, with the rest going to DFT. For LPMC, MLP retains a large weight, but this decreases while the weight for MNL and DFT increases with distance.
@@
The final step in our analysis of results concerns the performance of the model averaging structure on the test data. Table 5 shows the LL on the test data, covering the entire range of distances, for the four sub-models and for model averaging. As in estimation, we see that the model averaging structure obtains the highest LL. What is notable is the contrast with the results on the estimation data, i.e. Table 3. For both datasets, the gains by model averaging over the best performing sub-model are larger on the test data than on the estimation data.
@@
To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and test. We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 6 out of 8 for the LPMC data. When we turn to the test data, model averaging performs best (or joint best) in 2 out of 10 segments for the DECISIONS data, and 5 out of 10 segments for the LPMC data. These per segment statistics go hand in hand with the best overall LL obtained by model averaging as reported in Table 4 and Table 5.
@@
We earlier raised two possible reasons for why the advantages of MA might be larger on the test data than on the validation data. In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments (out of 2-9) where the per observation LL increases and segments where it decreases. No clear pattern emerges in the comparison across models and segments. Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.2582 vs -0.2944 for XGB). For the LPMC data, a similar story emerges. Model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10. Again, overall, MA performs the best of all models when looking at segments 1 and 10 together (-0.5793 vs -0.6025 for DFT).
@@ 4 Conclusions
There are, of course, some opportunities for further refinement. The key one relates to the fact that, in applying our proposed model averaging approach, an analyst needs to define what distribution means. In our work, we have focused on the obvious example of trip distance, for which there were pragmatic reasons, but also given that trip distance embodies key information in relation to similarity between trips. A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data. Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time. This, however, then also necessitates care by the analyst in defining what constitutes within-distribution and out-of-distribution.
@@ Acknowledgements
Stephane Hess acknowledges the support of the Delft Excellence Fund, as well as the European Research Council through the advanced grant 101020940-SYNERGY. Sander van Cranenburgh acknowledges the support of the TU Delft AI Labs programme. The authors are thankful to Thomas Hancock, Tim Hillel and Georges Sfeir for comments on an earlier draft.

### v3

@@ 3.4 Model averaging: specification and estimation
Similar to the MLP models used for training on the datasets (see Section 3.3), the MLP used for model averaging was implemented in PyTorch Paszke et al. (2019). This implementation is required because the loss function (see Equation (2)) involves a custom specification that cannot be readily implemented in, for example, Scikit-learn. As explanatory features, we included three distance-based variables: the raw trip distance, its logarithmic transformation, and its squared term. While in principle the linear distance alone should suffice, we found that including the logarithmic and squared transformations resulted in a small but consistent improvement in predictive performance. Figure 2 shows the structure for our model averaging approach. As shown in the figure, the model averaging MLP only takes the trip distance features as input to predict the weights given to each sub-model. The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model-averaging MLP.
Hyperparameters were tuned using a grid search combined with five-fold cross-validation, where folds were defined at the individual (household_id) level to avoid leakage across observations from the same individual. The following parameter spaces were considered: batch size (50, 100, 250), one hidden layer with 10, 20, or 30 nodes, learning rate (0.01, 0.001, 0.0001), and L2 regularisation strength (0, 0.01, 0.05). For the DECISIONS dataset, the best-performing model has a hidden layer with 10 nodes, a batch size of 50, a learning rate of 0.01, and an L2 regularisation strength of 0. For the LPMC dataset, the best-performing model has a hidden layer with 10 nodes, a batch size of 250, a learning rate of 0.01, and an L2 regularisation strength of 0.
@@
Using the optimal hyperparameters, we repeated the training 100 times. In contrast with the training of the individual models, our objective is now high predictive performance. We therefore retained only the top 20 per cent of models [^3] This selection was based on the log-likelihood. For this, we used the full data, i.e. training and validation data, but excluded the test data., discarding those that became trapped in poor local maxima. This approach differs slightly from the one used for training the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity, while diversity is now captured by the different models entering the ensemble.
It should be noted that the data used to train the meta-model are the same observations used to train the base models (for distance segments 3–8). Hastie et al. (2009) argue that using the same data for training both the base models and the meta-model may lead the meta-model to assign disproportionately high weights to more complex models, as their in-sample predictions can appear overly optimistic. To mitigate this issue, they propose a stacking procedure in which the meta-model is trained on predictions generated from data not used in training the base models (often obtained via cross-validated predictions). To further facilitate comparison with the machine learning literature on stacked generalisation, Table 3 provides a mapping between the terminology used in this paper and the standard terminology used in that literature.
While we recognise the concern about using the same for both the base models and meta-model, we deliberately refrained from implementing a stacking procedure in the main analysis in order to keep the modelling pipeline tractable. The overall framework already involves a large number of models, data partitions, hyperparameter tunings, and aggregation steps, and introducing a full cross-validated stacking procedure would substantially increase the computational and methodological complexity. Nevertheless, to assess whether this modelling choice affects our results, we conducted an additional analysis on the larger LPMC dataset in which the meta-model is trained on predictions generated from data not used to train the base models. The results of this analysis, reported in Appendix A, indicate that the impact on predictive performance is negligible, suggesting that the simpler procedure used in the main analysis does not materially affect the conclusions, at least in our application.
@@
Table 3: Terminology mapping between this study and the stacking literature
Current paper Stacking literature
Sub-model / individual model Base learner (level-0 model)
Model averaging model Meta-learner (level-1 model)
Model averaging weights Combination weights / stacking weights
Predicted probabilities (per model) Base learner predictions
Ensemble / model averaging approach Stacked generalisation / stacking
Distance-based features Meta-features
@@ 3.5.1 Performance of individual sub-models
Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 4 apply in each of the 6 distance segments used in estimation (i.e. segments 3-8). Across the distance segments, there are no cases where XGB outperforms MLP in the DECISIONS data, or MLP outperforms XGB in LPMC (which would reverse the overall trend), but differences change in size. Meanwhile, the overall trend between DFT and NL/MNL is reversed in some cases. These findings already show that different models perform differently well across trip distances. Interestingly, while the LPMC results suggest that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
@@
We next turn to the test results, shown in the right two graphs in Figure 3. We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data. For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for all models, with an average gain in per observation LL of $0.096$ for NL, $0.095$ for DFT, $0.062$ for MLP, and $0.065$ for XGB. The same happens for LPMC models, with improvements by $0.12$ for NL and DFT, $0.11$ for MLP and $0.081$ for XGB. Overall, these results do not suggest any overfitting, at least within the distribution of distances covered by the estimation data.
The findings for segments 2 and 9 on the test data largely confirm those we saw in the two graphs on the left in Figure 3, except for MLP and XGB performing more poorly in the test data for segment 9 for the DECISIONS data. We finally turn to segments 1 and 10. From a behavioural perspective, we see that for the DECISIONS data, mode choice on very short tips is much easier to predict. For short trips in the LPMC dataset, there is a clear advantage for the behavioural models. For the longest trips (segment 10), the trend is less clear. In the DECISIONS data, MLP and XGB perform comparatively well. In the LPMC data, both data-driven models again do comparatively well, while MNL suffers a big drop.
@@ 3.5.2 Model averaging results
For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e. distances on which the individual sub-models were estimated (segments 3-8). In the DECISIONS data, the weight of XGB increases on shorter trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP for DECISIONS. For LPMC, all the weight goes to XGB. These findings are in line with the pattern of prediction performance obtained for individual models, as shown on the left-hand side in Figure 3.
Once we go outside the $D_{train,sub}$ interval, the pattern changes. In both datasets, we observe that for trips with distances below $d_{a,sub}$ (i.e. segments 1&2), the weight for behavioural models increases. In the DECISIONS data, only a small share goes to NL and DFT. On the other hand, in LPMC, MNL obtains over 80% of the weight. For longer trips, i.e. above $d_{b,sub}$, NL gets the vast majority of the weight in the DECISIONS data, with DFT getting more than MLP or XGB. For LPMC, MLP retains a large weight, but this decreases while the weight for MNL and DFT increases with distance.
@@
The final step in our analysis of results concerns the performance of the model averaging structure on the test data. Table 6 shows the LL on the test data, covering the entire range of distances, for the four sub-models and for model averaging. As in estimation, we see that the model averaging structure obtains the highest LL. What is notable is the contrast with the results on the estimation data, i.e. Table 4. For both datasets, the gains by model averaging over the best-performing sub-model are larger on the test data than on the estimation data.
@@
To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and test. We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 6 out of 8 estimation segments for the DECISIONS data, and 7 out of 8 for the LPMC data. When we turn to the test data, model averaging performs best (or joint best) in 4 out of 10 segments for the DECISIONS data, and 5 out of 10 segments for the LPMC data. These per-segment statistics go hand in hand with the best overall LL obtained by model averaging as reported in Table 5 and Table 6.
@@
We earlier raised two possible reasons why the advantages of MA might be larger on the test data than on the validation data. In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments (out of 2-9) where the per observation LL increases and segments where it decreases. No clear pattern emerges in the comparison across models and segments. Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is higher than for any of the other models (-0.248 vs -0.252 for XGB). For the LPMC data, a similar story emerges. Model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10. Again, overall, MA performs the best of all models when looking at segments 1 and 10 together (-0.577 vs -0.6025 for DFT).
@@ 4 Conclusions
While the analysis focuses on predictive performance, the observed differences across model classes can be interpreted in light of their underlying structure. In particular, models with a stronger behavioural underpinning may exhibit more stable performance in out-of-distribution settings because their parameters are constrained by behavioural theory and regularity conditions. This can limit the extent to which such models capture spurious patterns or context-specific correlations present in the estimation data. In contrast, more flexible, data-driven models may achieve superior in-sample fit by capturing patterns that do not generalise, but this flexibility can lead to reduced robustness when extrapolating beyond the observed data. We emphasise that this interpretation remains tentative and is not directly tested in this study, but it provides a plausible explanation for the observed performance differences and aligns with broader discussions on the trade-off between flexibility and generalisation.
Our work relies on multiple separate models, and this clearly leads to increased complexity as well as doing nothing to address the black box nature of machine learning approaches. A contrasting approach is to make the trade-off between fit, generalisation, and interpretability more explicit at the estimation stage. Recent work has started to do so by proposing multi-objective estimation frameworks that jointly consider in-sample fit and out-of-sample predictive accuracy, thereby making it possible to explore Pareto-efficient specifications rather than selecting models on fit alone (Beeramoole et al., 2024). Related work has also shown that machine learning methods can be made more behaviourally credible by embedding behavioural regularity or interpretability constraints, which may help improve robustness when moving beyond the estimation domain (Feng et al., 2024). At the same time, the recent benchmarking literature makes clear that choice modellers are right to take the predictive strengths of machine learning seriously: across a very large empirical comparison, many ML models outperform traditional discrete choice models in prediction accuracy, even while important questions remain around transferability and uncertainty (Wang et al., 2024). It would therefore be unwise to disregard the strong in-sample performance that data-driven methods can deliver. Instead, our model averaging approach is motivated precisely by the idea of combining the strengths of both traditions: the behavioural interpretability and theoretical discipline of choice models, and the empirical flexibility and predictive power of machine learning methods.
There are, of course, some opportunities for further refinement. The key one relates to the fact that, in applying our proposed model averaging approach, an analyst needs to define what distribution means. In our work, we have focused on the obvious example of trip distance, for which there were pragmatic reasons, but also given that trip distance embodies key information in relation to similarity between trips. A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes besides trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data. Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time. This, however, then also necessitates care by the analyst in defining what constitutes within-distribution and out-of-distribution.
@@ Acknowledgements
Stephane Hess acknowledges the support of the Delft Excellence Fund, as well as the European Research Council through the advanced grant 101020940-SYNERGY. Sander van Cranenburgh acknowledges the support of the TU Delft AI Labs programme. The authors are thankful to Thomas Hancock, Tim Hillel and Georges Sfeir for comments on an earlier draft. The comments of three anonymous referees have helped us to substantially improve the paper.
Appendix A Cross-validated stacking LPMC
To assess whether the absence of a separate validation set for model averaging affects the results, we conducted an additional analysis using the larger LPMC dataset in which a dedicated split is introduced.
First, the base models (level-0 models) are trained exclusively on the estimation data. These trained models are then applied to a separate validation set that was not used during the training of the base models. The predictions generated by the base models on this validation set form the inputs for the model averaging stage.
Next, the model averaging MLP (level-1 model) is trained using this validation data. To facilitate training and prevent overfitting, the validation set is further split into a val_train and val_test subset. The val_train subset is used to estimate the parameters of the model averaging MLP, while the val_test subset is used to monitor predictive performance during training. Training is stopped early when performance on the val_test subset deteriorates, implementing an early stopping criterion.
This procedure ensures that the level-1 model is estimated using predictions obtained from data that were not used to train the base models, thereby providing a stricter separation between training stages. While this procedure is not identical to cross-validated stacking as typically implemented in the stacking literature, it closely mirrors the workflow used in the main analysis. As such, it provides a meaningful assessment of whether the absence of a dedicated validation set for model averaging affects the results reported in the main body of the paper.
The resulting splitting is illustrated in Figure 6.
Figure 6: Data splitting with validation
Below, we present the main results to facilitate comparison with those reported in the paper. The fit of the model averaging approach on the test data (covering distance segments 1-10) remains largely unchanged, with a log-likelihood of −11,204.87, vs. -11,115.43 in the paper, implying a drop in LL by $0.8\%$. More importantly, Figure 7 shows broadly the same mixing pattern as a function of distance. One difference is that the allocation of weights across models is somewhat sharper and less gradual. In particular, within distance segments 3–8 virtually all weight is assigned to XGB, whereas outside this range XGB receives almost no probability mass. In segment 9, MLP attains the largest share of the probability, while segments 1 and 10 exhibit a mixture of MNL, DFT, and MLP.
Figure 7: Heatmap of model performance across distance segments.
A comparison of Figure 8 with the results reported in the paper reveals a largely similar pattern. The numerical values change only marginally. The only noteworthy difference is a modest increase in the mean log-likelihood achieved by model averaging in distance segment 1, which rises to 0.726-0.726 (vs. 0.658-0.658 in the paper). The slightly sharper allocation may be related to the revised estimation procedure, where the level-0 models and the meta-model are trained on separate data partitions following Hastie et al. (2009). Training the meta-model on predictions that reflect out-of-sample performance can affect the allocation of weights across models. However, given the very modest changes observed, the results suggest that not using a separate validation set for model averaging has little impact on our overall findings.
Figure 8: Heatmap of model performance across distance segments.